Q.write the product .
1. (x+a)(x-b)(x-c)
2. (x+a+b)(x+c-d)

Answer by Alan3354(69443)

1. (x+a)(x-b)(x-c)
Multiply any 2
(x+a)(x-b) = x*(x-b) + a*(x-b) = x^2 - bx + ax - ab or (x^2 + (a-b)x - ab)
Then multiply that by (x-c)
=================
2. (x+a+b)(x+c-d)